OpenStudy (anonymous):

You have two beakers on your lab table. Beaker 1 contains 32.07 g of sulfur and Beaker 2 contains 74.92 g of arsenic (As). Which beaker contains the greatest number of atoms? Choose the best answer.

OpenStudy (noelgreco):

The only way to compare is to convert to moles.

OpenStudy (joannablackwelder):

And then to number of atoms. :)

OpenStudy (abmon98):

1)Number of moles=Mass(g)/Molar Mass(g/mol) 2) Avogadro's number is the number of particles found in one mole of a substance. the determined value is approximately 6.022*10^23 particles per mole use a ratio when you obtains each of the two beakers number of moles

OpenStudy (anonymous):

(32.07 g S) / (32.0655 g S/mol) = 1.0001 mol S (74.92 g As) / (74.92160 g As/mol) = 0.99998 mol As The number of atoms is proportional to the number of moles, so there is a slightly greater number of S atoms. But as a practical matter, the difference between the two lies outside the number of significant digits given in the question, so really the two beakers should be judged to have the same number of atoms, within the accuracy given in the question.